# Extrapolation for bilinear compact operators in the variable exponent setting

### v1

@@ Abstract.
We establish extrapolation of compactness for bilinear operators in the scale of weighted variable exponent Lebesgue spaces. First, we prove an abstract principle relying on the Cobos–Fernández–Cabrera–Martínez theorem. Then, as an application we deduce new compactness results for the commutators of bilinear $\omega$-Calderón–Zygmund operators and bilinear fractional integrals acting on weighted variable exponent Lebesgue spaces. Our work extends and unifies among others earlier works of the second named author together with Hytönen as well as Oikari.
@@ 1. Introduction
Such boundedness results and quantitative versions of them are often proved more easily for $p=2$, among other reasons, because $L^{2}(w)$ is a Hilbert space. To extend them to the whole range $1<p<\infty$, one needs an extrapolation result. In this case, the relevant method was introduced by Rubio de Francia [34]. Several works followed further developing this method, until Duoandikoetxea [14] proved a sharp version of it. We refer to the book [9] for a thorough historical overview as well as detailed proofs and generalizations.
There exist several directions for studying more general problems than the above mentioned boundedness. One consists in considering multilinear Calderón–Zygmund operators. Lerner, Ombrosi, Pérez, Torres and Trujillo-González [23] showed that if $T$ is a $m$-linear Calderón–Zygmund operator on $\mathbb{R}^{n}$, $\vec{p}=(p_{1},\ldots,p_{m})$ is a $m$-tuple of exponents from $(1,\infty)$, and $\vec{w}=(w_{1},\ldots,w_{m})$ is a $m$-tuple of weights such that
@@
then $T$ maps $L^{p_{1}}(w_{1})\times\ldots\times L^{p_{m}}(w_{m})$ into $L^{p}(\prod_{j=1}^{m}w_{j})$ boundedly. Extrapolation results of such boundedness have played here also a key role in the further development of the theory, see [32].
Another direction consists in considering weighted Lebesgue spaces $L^{p(\cdot)}(w)$ with a variable exponent $p(\cdot)$, that means $p(\cdot)$ is now a not necessarily constant function $p(\cdot):\mathbb{R}^{n}\to[0,\infty]$. For the precise definition of these spaces we refer to Section 2. An exhaustive presentation of the unweighted variable exponent Lebesgue spaces $L^{p(\cdot)}(\mathbb{R}^{n})$ can be found in [12, 13]. Cruz-Uribe, Fiorenza and Neugebauer [8] showed that for every $p(\cdot)\in\mathscr{P}\cap LH$ (see Section 2 for the notation), the Hardy–Littlewood maximal operator is bounded on $L^{p(\cdot)}(w)$ if any only if
@@
is finite. This was extended in [33] to the bilinear setting, where moreover analogous boundedness results for bilinear singular integrals were proved. Also, there is by now a rich extrapolation theory for variable exponent spaces both in the linear and the multilinear setting, see [11, 4, 31]. However, it is still unknown if one can prove the exact linear and multilinear analogue of Rubio de Francia’s extrapolation of boundedness on weighted Lebesgue spaces with variable exponents (see also [22, Remark 1.7]).
Over the last few years, the extension of extrapolation results to compact operators has drawn the attention of many authors. In particular, the first work that dealt with this mathematical problem in the linear setting was due to Hytönen and the second named author [17]. Further extensions of this result in the setting of weighted Morrey spaces, two weight problems, Banach function spaces and one sided situations can be found in [21], [24], [26] (see also [22] for an alternative proof in the context of weighted Lebesgue spaces with variable exponents) and [29], respectively. Moreover, first Cao, Olivo and Yabuta [5] established extrapolation results for multilinear compact operators (see also [18]).
@@
Motivated by the previously mentioned works, in this paper we establish the following abstract diagonal and off-diagonal extrapolation theorem for bilinear compact operators in the setting of weighted variable exponent Lebesgue spaces (see Section 2 for detailed definitions).
@@
Let $T$ be a bilinear operator such that there exist constants $t\in[1,\infty)$ and $\gamma\in\left[0,\frac{2}{t}\right)$ satisfying both of the following two assumptions.
@@
    For all $\vec{p}_{0}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{2}$, $q_{0}(\cdot)\in\mathcal{E}^{0}_{(1,\infty)}$ with
@@
    $\frac{1}{p_{0}(\cdot)}=\frac{1}{p_{0,1}(\cdot)}+\frac{1}{p_{0,2}(\cdot)},\quad\frac{1}{p_{0}(\cdot)}-\frac{1}{q_{0}(\cdot)}=\gamma,$
    the following holds. Whenever $\vec{w}_{0}$ is a $2$-tuple of weights such that $\vec{w}_{0}\in\mathcal{A}^{\gamma}_{\vec{p}_{0}(\cdot),q_{0}(\cdot)}$ and $\vec{w}_{0}^{t}\in\mathcal{A}^{t\gamma}_{\frac{\vec{p}_{0}(\cdot)}{t},\frac{q_{0}(\cdot)}{t}}$, we have
@@
    There exist some fixed $2$-tuple of variable exponents $\vec{p}_{1}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{2}$, a fixed variable exponent $q_{1}(\cdot)\in\mathcal{E}^{0}_{(1,\infty)}$ with
@@
    $\frac{1}{p_{1}(\cdot)}=\frac{1}{p_{1,1}(\cdot)}+\frac{1}{p_{1,2}(\cdot)},\quad\frac{1}{p_{1}(\cdot)}-\frac{1}{q_{1}(\cdot)}=\gamma,$
    and a fixed $2$-tuple of weights $\vec{w}_{1}$ with $\vec{w}_{1}\in\mathcal{A}^{\gamma}_{\vec{p}_{1}(\cdot),q_{1}(\cdot)}$ and $\vec{w}_{1}^{t}\in\mathcal{A}^{t\gamma}_{\frac{\vec{p}_{1}(\cdot)}{t},\frac{q_{1}(\cdot)}{t}}$, such that
@@
Then: For all $\vec{p}(\cdot)\in(\mathcal{E}^{0}_{(1,\infty)})^{2}$, $q(\cdot)\in\mathcal{E}^{0}_{(1,\infty)}$, such that
@@
$\frac{1}{p(\cdot)}=\frac{1}{p_{1}(\cdot)}+\frac{1}{p_{2}(\cdot)},\quad\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}=\gamma,$
and for all $2$-tuples of weights $\vec{w}$ with $\vec{w}\in\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}$ and $\vec{w}^{t}\in\mathcal{A}^{t\gamma}_{\frac{\vec{p}(\cdot)}{t},\frac{q(\cdot)}{t}}$, we have
@@
As applications, we obtain new compactness results for the commutators of bilinear $\omega$-Calderón–Zygmund operators and bilinear fractional integrals.
@@
By assuming $t=1$ and choosing all the variable exponents to be constant ones in the assumptions and conclusion of Theorem 1.1, we can extend or recover several results of [5, 18, 22, 26].
The remainder of this paper is organized as follows. In Section 2, we recall the definition of bilinear compact operators and state several definitions about unweighted and weighted variable Lebesgue spaces. In Section 3, we gather some auxiliary lemmas that will aid us in proving our main result. It is worth mentioning that the proof of one of these lemmas, namely, the direction “if” of Lemma 3.4 is new. Section 4 is devoted to the proof of two key factorization results in the setting of variable multilinear Muckenhoupt weights (see Lemmas 4.1 and 4.2). In Section 5, we recall several known facts about complex interpolation for bilinear compact operators. Among them is the abstract result of Cobos–Fernández-Cabrera–Martínez [7, Theorem 3.2]. Lastly, in Sections 6 and 7, we provide the proof of Theorem 1.1 and its applications, respectively.
@@ 2. Preliminaries
Throughout the paper, we mostly consider bilinear compact operators. Hence, in this section we recall this definition for the reader’s convenience. Given the normed spaces $X,Y$ and $Z$, a bilinear operator $T:X\times Y\rightarrow Z$ is said to be compact if the set $\{T(x,y):\;\|x\|\leq 1,\;\;\|y\|\leq 1\}$ is precompact in $Z$. Some other equivalent formulations of this definition were given in [2, Proposition 1].
@@ 2.A. Variable Lebesgue spaces
In this subsection, we recall some basic definitions of variable Lebesgue spaces. For more information, we refer the reader to [12] and [13].
@@ 2.A.1. Notation for variable exponent functions
$\mathcal{P}_{0}:=\big\{\text{all Lebesgue measurable functions }p(\cdot):\mathbb{R}^{n}\to(0,\infty]\big\},$
$\mathscr{P}_{0}:=\{p(\cdot)\in\mathcal{P}_{0}:~p_{+}<\infty\},$
$\mathcal{P}:=\{p(\cdot)\in\mathcal{P}_{0}:~p(x)\geq 1,\text{ for all }x\in\mathbb{R}^{n}\},$
@@
$\mathscr{P}:=\{p(\cdot)\in\mathcal{P}:~1<p_{-}\leq p_{+}<\infty\}.$
@@ 2.A.2. Luxemburg norm of variable Lebesgue spaces
Given $p(\cdot)\in\mathcal{P}_{0}$, define the modular associated with $p(\cdot)$ by
(2.1) $\rho_{p(\cdot)}(f):=\int_{\mathbb{R}^{n}\setminus\Omega_{\infty}}|f(x)|^{p(x)}\,\mathrm{d}x+\|f\|_{L^{\infty}(\Omega_{\infty})},$
where $\Omega_{\infty}:=\{x\in\mathbb{R}^{n}\;:\;p(x)=\infty\}$. The variable Lebesgue space $L^{p(\cdot)}(\mathbb{R}^{n})$ consists of all measurable functions $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ such that
@@
In what follows, we always assume that $p_{+}<\infty$ and hence the latter term in (2.1) is zero.
@@ 2.A.4. Log-Hölder continuity
    it is locally log-Hölder continuous, denoted by $p(\cdot)\in LH_{0},$ if there exists $C_{0}>0$ such that for all $x,y\in\mathbb{R}^{n}$, whenever $|x-y|<\tfrac{1}{2}$, then
@@
    it is log-Hölder continuous at infinity, denoted by $p(\cdot)\in LH_{\infty},$ if there exist $p_{\infty}\in\mathbb{R}$ and $C_{\infty}>0$ such that for all $x\in\mathbb{R}^{n}$,
@@
We define $LH:=LH_{0}\cap LH_{\infty}.$
@@ 2.B. Variable exponent Muckenhoupt characteristics
In what follows, we work directly in the off-diagonal setting, namely, we consider $\gamma\in(0,1)$ and not only $\gamma=0$. Let $\gamma\in[0,1)$ be fixed and denote
(2.3) $\mathcal{E}_{(1,\infty)}^{0}$ $:=LH\cap\mathscr{P},$
$\mathcal{E}^{\gamma}_{(1,\infty)}$ $:=\left\{\big(p(\cdot),q(\cdot)\big)\in\mathcal{E}_{(1,\infty)}^{0}\times\mathcal{E}_{(1,\infty)}^{0}:1/p(\cdot)-1/q(\cdot)=\gamma\right\}.$
The appropriate class of weights is as follows.
2.4 Definition ([11], Definition 2.10).
Let $\gamma\in[0,1)$ and $p(\cdot),q(\cdot)\in\mathcal{P}$ be such that $1/p(\cdot)-1/q(\cdot)=\gamma.$ Then, a weight $w$ belongs to the class $\mathcal{A}_{p(\cdot),q(\cdot)}^{\gamma},$ if
$w]_{\mathcal{A}_{p(\cdot),q(\cdot)}^{\gamma}}:=\sup_{Q}|Q|^{\gamma-1}\|w\chi_{Q}\|_{q(\cdot)}\|w^{-1}\chi_{Q}\|_{p^{\prime}(\cdot)}<\infty,$
where the supremum is taken over all cubes.
When $\gamma=0$ we understand that $p(\cdot)\in\mathcal{E}_{(1,\infty)}^{0}$ is equivalent to $(p(\cdot),p(\cdot))\in\mathcal{E}_{(1,\infty)}^{0}$ and $\mathcal{A}_{p(\cdot),q(\cdot)}^{0}=\mathcal{A}_{p(\cdot)}$.
@@
Given variable exponents $p_{1}(\cdot),\dots,p_{m}(\cdot)$, $p_{0,1}(\cdot),\dots,p_{0,m}(\cdot)$ and $p_{1,1}(\cdot),\dots,\\ p_{1,m}(\cdot)$ we define the vectors
$\vec{p}(\cdot):=(p_{1}(\cdot),\dots,p_{m}(\cdot)),\quad\vec{p}_{0}(\cdot):=(p_{0,1}(\cdot),\dots,p_{0,m}(\cdot))\quad\text{and}\quad\vec{p}_{1}(\cdot):=(p_{1,1}(\cdot),\dots,p_{1,m}(\cdot)).$

### v2

@@ Abstract.
We establish extrapolation of compactness for bilinear operators in the scale of weighted variable exponent Lebesgue spaces. First, we prove an abstract principle relying on the Cobos–Fernández-Cabrera–Martínez theorem. Then, as an application we deduce new compactness results for the commutators of bilinear $\omega$-Calderón–Zygmund operators, bilinear fractional integrals and bilinear Fourier multipliers acting on weighted variable exponent Lebesgue spaces. Our work extends and unifies among others earlier works of the second named author together with Hytönen as well as Oikari.
@@ 1. Introduction
Such boundedness results and quantitative versions of them are often proved more easily for $p=2$, among other reasons, because $L^{2}(w)$ is a Hilbert space. To extend them to the whole range $1<p<\infty$, one needs an extrapolation result. In this case, the relevant method was introduced by Rubio de Francia [undefaj]. Several works followed further developing this method, until Duoandikoetxea [undefm] proved a sharp version of it. We refer to the book [undefh] for a thorough historical overview as well as detailed proofs and generalizations.
There exist several directions for studying more general problems than the above mentioned boundedness. One consists in considering multilinear Calderón–Zygmund operators. Lerner, Ombrosi, Pérez, Torres and Trujillo-González [undefw] showed that if $T$ is a $m$-linear Calderón–Zygmund operator on $\mathbb{R}^{n}$, $\vec{p}=(p_{1},\ldots,p_{m})$ is a $m$-tuple of exponents from $(1,\infty)$, and $\vec{w}=(w_{1},\ldots,w_{m})$ is a $m$-tuple of weights such that
@@
then $T$ maps $L^{p_{1}}(w_{1})\times\cdots\times L^{p_{m}}(w_{m})$ into $L^{p}(\prod_{j=1}^{m}w_{j})$ boundedly. Extrapolation results of such boundedness have played here also a key role in the further development of the theory, see [undefah].
Another direction consists in considering weighted Lebesgue spaces $L^{p(\cdot)}(w)$ with a variable exponent $p(\cdot)$, that means $p(\cdot)$ is now a not necessarily constant function $p(\cdot):\mathbb{R}^{n}\to[0,\infty]$. For the precise definition of these spaces we refer to Section 2. An exhaustive presentation of the unweighted variable exponent Lebesgue spaces $L^{p(\cdot)}(\mathbb{R}^{n})$ can be found in [undefk, undefl]. Cruz-Uribe, Fiorenza and Neugebauer [undefg] showed that for every $p(\cdot)\in\mathscr{P}\cap\mathrm{LH}$ (see Section 2 for the notation), the Hardy–Littlewood maximal operator is bounded on $L^{p(\cdot)}(w)$ if any only if
@@
is finite. This was extended in [undefai] to the bilinear setting, where moreover analogous boundedness results for bilinear singular integrals were proved. Also, there is by now a rich extrapolation theory for variable exponent spaces both in the linear and the multilinear setting, see [undefj, undefc, undefag]. However, it is still unknown if one can prove the exact linear and multilinear analogue of Rubio de Francia’s extrapolation of boundedness on weighted Lebesgue spaces with variable exponents (see also [undefv, Remark 1.7]).
Over the last few years, the extension of extrapolation results to compact operators has drawn the attention of many authors. In particular, the first work that dealt with this mathematical problem in the linear setting was due to Hytönen and the second named author [undefq]. Further extensions of this result in the setting of weighted Morrey spaces, two weight problems, Banach function spaces and one sided situations can be found in [undefu], [undefy], [undefaa] (see also [undefv] for an alternative proof in the context of weighted Lebesgue spaces with variable exponents) and [undefad], respectively. Moreover, first Cao, Olivo and Yabuta [undefd] established extrapolation results for multilinear compact operators (see also [undefr]).
@@
Motivated by the previously mentioned works, in this paper we establish the following abstract diagonal and off-diagonal extrapolation, full range and limited range theorems for bilinear compact operators in the setting of weighted variable exponent Lebesgue spaces (see Section 2 for detailed definitions and explanation of the notation).
@@
Let $T$ be a bilinear operator such that there exist fixed constants $t\in(0,\infty)$ and $\gamma\in\left[0,\infty\right)$ satisfying both of the following two assumptions.
@@
    For all proper $2$-admissible quadruples $(\vec{p}_{0}(\cdot),q_{0}(\cdot),\vec{1},\infty)$ with
    $(q_{0})_{-}>1,$
@@
    $\frac{1}{p_{0}(\cdot)}-\frac{1}{q_{0}(\cdot)}=\gamma,$
    the following holds. Whenever $\vec{w}_{0}$ is a $2$-tuple of weights such that $\vec{w}_{0}\in\mathcal{A}_{\vec{p}_{0}(\cdot),q_{0}(\cdot)}$ and $\vec{w}_{0}^{t}\in\mathcal{A}_{\frac{\vec{p}_{0}(\cdot)}{t},\frac{q_{0}(\cdot)}{t}}$, we have
@@
    There exist some fixed proper $2$-admissible quadruple $(\vec{p}_{1}(\cdot),q_{1}(\cdot),\vec{1},\infty)$ with
    $(q_{1})_{-}>1,$
@@
    $\frac{1}{p_{1}(\cdot)}-\frac{1}{q_{1}(\cdot)}=\gamma,$
    and a fixed $2$-tuple of weights $\vec{w}_{1}$ with $\vec{w}_{1}\in\mathcal{A}_{\vec{p}_{1}(\cdot),q_{1}(\cdot)}$ and $\vec{w}_{1}^{t}\in\mathcal{A}_{\frac{\vec{p}_{1}(\cdot)}{t},\frac{q_{1}(\cdot)}{t}}$, such that
@@
Then: For all proper $2$-admissible quadruples $(\vec{p}(\cdot),q(\cdot),\vec{1},\infty)$ such that
$q_{-}>1,$
@@
$\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}=\gamma,$
and for all $2$-tuples of weights $\vec{w}$ with $\vec{w}\in\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}$ and $\vec{w}^{t}\in\mathcal{A}_{\frac{\vec{p}(\cdot)}{t},\frac{q(\cdot)}{t}}$, we have
@@
1.2 Theorem.
Let $T$ be a bilinear operator such that there exist a fixed constant $\gamma\in\left[0,\infty\right)$, a fixed $m$-tuple $\vec{r}\in[1,\infty)^{m}$ and a fixed constant $s\in(0,\infty]$ satisfying both of the following two assumptions.
  1. (1)
    For all proper $2$-admissible quadruples $(\vec{p}_{0}(\cdot),q_{0}(\cdot),\vec{r},s)$ with
    $(q_{0})_{-}>1$
    and
    $\frac{1}{p_{0}(\cdot)}-\frac{1}{q_{0}(\cdot)}=\gamma,$
    the following holds. Whenever $\vec{w}_{0}$ is a $2$-tuple of weights such that $\vec{w}_{0}\in\mathcal{A}_{(\vec{p}_{0}(\cdot),q_{0}(\cdot)),(\vec{r},s)}$, we have
    boundedly.
  2. (2)
    There exist some fixed proper $2$-admissible quadruple $(\vec{p}_{1}(\cdot),q_{1}(\cdot),\vec{r},s)$ with
    $(q_{1})_{-}>1$
    and
    $\frac{1}{p_{1}(\cdot)}-\frac{1}{q_{1}(\cdot)}=\gamma,$
    and a fixed $2$-tuple of weights $\vec{w}_{1}$ with $\vec{w}_{1}\in\mathcal{A}_{(\vec{p}_{1}(\cdot),q_{1}(\cdot)),(\vec{r},s)}$, such that
    compactly.
Then: For all proper $2$-admissible quadruples $(\vec{p}(\cdot),q(\cdot),\vec{r},s)$ such that
$q_{-}>1$
and
$\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}=\gamma,$
and for all $2$-tuples of weights $\vec{w}$ with $\vec{w}\in\mathcal{A}_{(\vec{p}(\cdot),q(\cdot)),(\vec{r},s)}$, we have
compactly.
As applications of Theorem 1.1, we obtain new compactness results for the commutators of bilinear $\omega$-Calderón–Zygmund operators and bilinear fractional integrals (see Subsections 7.C and 7.D). Moreover, as an application of Theorem 1.2, we obtain new compactness results for the commutators of bilinear Fourier multipliers (see Subsection 7.E).
@@
By assuming $t=1$ and choosing all the variable exponents to be constant in the assumptions and conclusions of Theorems 1.1 and 1.2, we can extend or recover several results of [undefd, undefr, undefaa] in the bilinear case.
The remainder of this paper is organized as follows. In Section 2, we recall the definition of bilinear compact operators and state several definitions about unweighted and weighted variable Lebesgue spaces. In Section 3, we gather some auxiliary lemmata that will aid us in proving our main results. As far as we know, one of these lemmata, namely, Lemma 3.5 is new. Section 4 is devoted to the proof of two key factorization results in the setting of variable multilinear Muckenhoupt weights (see Lemmata 4.1 and 4.3). In Section 5, we recall several known facts about complex interpolation for bilinear compact operators. Among them is the abstract result of Cobos–Fernández-Cabrera–Martínez [undeff, Theorem 3.2]. Lastly, in Sections 6 and 7, we provide the proofs of Theorems 1.1 and 1.2 and their applications, respectively.
@@ 2. Preliminaries
Throughout the paper, we mostly consider bilinear compact operators. Hence, in this section we recall this definition for the reader’s convenience. Given the normed spaces $X,Y$ and $Z$, a bilinear operator $T:X\times Y\rightarrow Z$ is said to be compact if the set $\{T(x,y):\;\|x\|\leq 1,\;\;\|y\|\leq 1\}$ is precompact in $Z$. Some other equivalent formulations of this definition were given in [undefa, Proposition 1].
@@ 2.A. Variable Lebesgue spaces
In this subsection, we recall some basic definitions of variable Lebesgue spaces. For more information, we refer the reader to [undefk] and [undefl].
@@ 2.A.1. Notation for variable exponent functions
$\mathscr{P}_{0}:=\{p(\cdot):~0<p_{-}\leq p_{+}<\infty\},$
$\mathcal{P}:=\{p(\cdot):~1\leq p_{-}\leq p_{+}<\infty\},$
@@
$\mathscr{P}:=\{p(\cdot):~1<p_{-}\leq p_{+}<\infty\}.$
@@ 2.A.2. Luxemburg norm of variable Lebesgue spaces
Given $p(\cdot)\in\mathscr{P}_{0}$, define the modular associated with $p(\cdot)$ by
(2.1) $\rho_{p(\cdot)}(f):=\int_{\mathbb{R}^{n}}|f(x)|^{p(x)}\,\mathrm{d}x.$
The variable Lebesgue space $L^{p(\cdot)}(\mathbb{R}^{n})$ consists of all measurable functions $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ such that
@@ 2.A.4. Log-Hölder continuity
    it is locally log-Hölder continuous, denoted by $p(\cdot)\in\mathrm{LH}_{0},$ if there exists $C_{0}>0$ such that for all $x,y\in\mathbb{R}^{n}$, whenever $|x-y|<\tfrac{1}{2}$, then
@@
    it is log-Hölder continuous at infinity, denoted by $p(\cdot)\in\mathrm{LH}_{\infty},$ if there exist $p_{\infty}\in\mathbb{R}$ and $C_{\infty}>0$ such that for all $x\in\mathbb{R}^{n}$,
@@
We define $\mathrm{LH}:=\mathrm{LH}_{0}\cap\mathrm{LH}_{\infty}.$
@@ 2.B. Variable exponent Muckenhoupt characteristics
Here we define the main classes of weights that we will be considering in this paper.
@@
  1. (1)
    A quadruple $(\vec{p}(\cdot),q(\cdot),\vec{r},s)$ is said to be $m$-admissible if all of the following conditions are satisfied:
    • $\vec{r}=(r_{1},\ldots,r_{m})\in(0,\infty)^{m}$ is a $m$-tuple of constants.
    • $s\in(0,\infty]$ is a constant.
    • $\vec{p}(\cdot)=(p_{1}(\cdot),\ldots,p_{m}(\cdot))\in\mathscr{P}_{0}^{m}$ is a $m$-tuple of variable exponents with
      $r_{j}<(p_{j})_{-},\quad\forall j=1,\ldots,m.$
    • $q(\cdot)\in\mathscr{P}_{0}$ is a variable exponent with $q_{+}<s$.
    • There is a constant $\gamma\in[0,\infty)$ such that
      $\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}=\gamma,$
      where the variable exponent $p(\cdot)\in\mathscr{P}_{0}$ is defined via
      $\frac{1}{p(\cdot)}:=\sum_{j=1}^{m}\frac{1}{p_{j}(\cdot)}.$
    In this case, we define the constant $r\in(0,\infty)$ via
    $\frac{1}{r}:=\sum_{j=1}^{m}\frac{1}{r_{j}}.$
  2. (2)
    A $m$-admissible quadruple $(\vec{p}(\cdot),q(\cdot),\vec{r},s)$ is said to be proper if
    $p_{j}(\cdot)\in\mathrm{LH},\quad\forall j=1,\ldots,m$
    and
    $q(\cdot)\in\mathrm{LH}.$
    Observe that in this case, since $(p_{j})_{+}<\infty$, by [undefk, Proposition 2.3] we know that $p_{j}(\cdot)\in\mathrm{LH}$ if and only if $1/p_{j}(\cdot)\in\mathrm{LH}.$ Hence, $\frac{1}{p(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{j}(\cdot)}$ is a linear combinations of functions in $\mathrm{LH}$. Thus, combining this with the fact that $\mathrm{LH}$ is closed under linear combinations we conclude that $p(\cdot)\in\mathrm{LH}$.
For technical reasons and also for comparing with earlier definitions that were given in the literature, we observe the following.
2.4 Lemma.
Let $(\vec{p}(\cdot),q(\cdot),\vec{r},s)$ be a $m$-admissible quadruple. Set
$\rho_{1}:=r$
and define $\rho_{2}\in(0,\infty)$ via
$\frac{1}{\rho_{1}}-\frac{1}{\rho_{2}}:=\gamma.$
Moreover, set
$s_{2}:=s$
and define $s_{1}\in(0,\infty]$ via
$\frac{1}{s_{1}}-\frac{1}{s_{2}}:=\gamma.$
Then, there holds
$\rho_{1}<p_{-}\leq p_{+}<s_{1},$
and
$\rho_{2}<q_{-}\leq q_{+}<s_{2}.$
Moreover, $\rho_{1}\leq\rho_{2}$ and $s_{1}\leq s_{2}$. In particular, if $\rho_{1}\geq 1$, then also $\rho_{2}\geq 1$.
Proof.
The assertions of the last two sentences are obvious. Thus, we concentrate on proving the rest of the lemma.
First of all, observe that
$\frac{1}{r}=\sum_{i=1}^{m}\frac{1}{r_{j}}>\sum_{i=1}^{m}\frac{1}{(p_{j})_{-}}\geq\sum_{i=1}^{m}\frac{1}{p_{j}(\cdot)}=\frac{1}{p(\cdot)},$
so by taking the essential supremum we arrive at
$\rho_{1}=r<p_{-}.$
Moreover, we have
(2.5) $\frac{1}{p(\cdot)}=\frac{1}{q(\cdot)}+\gamma,$
therefore by taking the essential supremum we arrive at
$\frac{1}{p_{-}}=\frac{1}{q_{-}}+\gamma,$
so in particular
$\frac{1}{\rho_{1}}=\frac{1}{r}>\frac{1}{p_{-}}>\gamma,$
showing in particular that indeed we can define $\rho_{2}\in(0,\infty)$ as in the statement of the lemma. Furthermore, we have
$q_{-}=\frac{1}{\frac{1}{p_{-}}-\gamma}>\frac{1}{\frac{1}{\rho_{1}}-\gamma}=\rho_{2}.$
In addition, we have $q_{+}<s=s_{2}$ by definition. Since
$\frac{1}{s_{2}}+\gamma\geq 0,$
one can indeed define $s_{1}\in(0,\infty]$ as in the statement of the lemma. Finally, by taking the essential infimum of both sides of (2.5), we have
$\frac{1}{p_{+}}=\frac{1}{q_{+}}+\gamma.$
Thus, we compute
$p_{+}=\frac{1}{\gamma+\frac{1}{q_{+}}}<\frac{1}{\gamma+\frac{1}{s}}=s_{1},$
concluding the proof. ∎